sum of a number and its double is 9.Find the number.
Found 2 solutions by tommyt3rd, lenny460:
Answer by tommyt3rd(5050)

actually since 9 is such a small number we don't really need algebra: 3 works.
But anyways...
a number--x
its double---2x
their sum is 9 means:
x+2x=9
3x=9
(3x)/3=9/3
x=3

:)

Answer by lenny460(1073)

The number is 3.